INAF LBC+LUCIFER queue observing, 2010 April10-11

Observer: A. Fontana, K.Boutsia, R. Speziali
Telescope Operator: D.Gonzales Huerta
Support Astronomer: D. Thompson
Instrument: LBC

Summary

Clear night but not photometric.
30 minutes lost because of problems with the rotator
30 minutes lost because interface got blocked. We had to do : LBC kill - LBC restart
2.5h lost because of guiding star failure. The problem was due to the expired IDL licence. A. di Paola modified the script and we went on without AO utility

Data Summary

Project Objects Filter ExpTime Notes
Giallongo Q0933 r-SLOAN 1.5h some exposures are done without guiding
Ferraro n5466 B-bessel ~1.8h  
Ferraro n5466 r-SLOAN ~1.8h  
Calibration BD+26 Us, g, r   probably saturated

Details

01:46 - Bias
  • 5Bias_Bino_5
  • 5Bias_Bino_rot90 - 2 times because problems with the mirror
  • 5Bias-Bino_rot180
  • 5Bias-Bino_rot270

02:00 - Pointing for flats
  • Flat script in binocular mode STD_Usp and r-SLOAN (P.A 0 and P.A 180)

02:40 - Slew to Q0933 (Giallongo) - correct pointing near the field
  • pointing star WT10_305

02:49 - Slew again to Q0933 (Giallongo) - dofpia
  • Blue fails to converge, too big ficus corrections and exits
  • we apply z7=-3000, z4=+15000, z11=+1000
  • fail again
  • dofpia, /Blueonly -> it converges with values z7~-1000 and z4~+2000

03:22 - dofpia - binocular

03:26 - preset to Q0933 - start OB (0.5h)

04:19 - Pointing correction - PROBLEMS WITH POINTING
  • correct rotator center not written in the header - ptadjust - mirroradjust

04:56 - Slew to Q0933 - dofpia

04:59 - Start integration Q0933
  • interface got blocked after 4th exposure
  • we had to restart interface by doing : LBC kill - LBC restart

05:38 - dofpia - after recovery from LBC crash

05:43 - Trackers are giving errors - reboot LBC

05:55 - Preset to Q0933 - repeat OB
  • guiding seems correct

06:00 - Images look out of focus - Abort OB for collimation
  • dofpia

06:03 - Preset to Q0933 - repeat OB
  • NOT guiding - Abort OB
  • Still giving tracker errors - ERROR: RPC Client/Host communication failed retval:5 [src/trackers/trackers.c:1728]
  • reboot LBC

06:18 - dofpia

06:23 - Q0933_Usr-SLOAN_2.ob - repeated
  • guide star failed
  • Call Andrea Di Paola!!!

06:47 - Blue arm disabled - repeat Q0933 OB with Red only
  • guiding failed again but we do not interrupt OB

07:54 - Andrea found that the problem is due to the EXPIRATION OF THE IDL LICENSE! The IDL routine called by LBCIARUN.EXE was opening a pop up window that hung all the processes. So Andrea changed the name of the LBCIARUN.EXE not to call IDL anymore. It introduced a non critical error but allowing us to go on even if without the AO utility. Start an Q0933 OB with the Blue arm only for test. * LBC kill - LBC start before the test OB

08:18 - Preset to Ferraro n5466_f1_1 - dofpia

08:25 - Exposure - n5466_f1_1

08:42 - Exposure - n5466_f1_2

08:57 - Exposure - n5466_f1_3

09:05 - Abort OB because there is condensation on the RED camera.
  • close shutter for 5 minutes and wait for air flow to clear the camera
  • Roberto goes to check camera in the dome

09:35 - Test image at the horizon - everything seem to be working
  • we changed exposure time in the Ferraro OBs - 200x4 instead of 400x2 (RED) / 410x2 instead of 400x2

09:45 - dofpia

09:50 - repeat n5466_f1_3 OB
  • OB executed with new exposure times

10:08 - Exposure - n5466_f1_4
  • OB executed with new exposure times

10:25 - dofpia

10:31 - Exposure - n5466_f1_5
  • OB executed with new exposure times

10:48 - Exposure - n5466_f2_1
  • OB executed with new exposure times

11:05 - dofpia

11:09 - Exposure - n5466_f2_2
  • OB executed with new exposure times

11:26 - Exposure - n5466_f2_3
  • OB executed with new exposure times

11:44 - dofpia

11:49 - Exposure - n5466_f2_4
  • OB executed with new exposure times
  • OB has to be repeated - out of specifics

12:05 - Slew to Photometric Standard - BD+26Usg_r.ob

12:13 - Slew to Photometric Standard - G138_31_UsB_r.ob
  • aborted because saturated
  • both standards seem to be saturated

12:27 - Sky Flats in B-bessel and r-SLOAN
  • counts are somewhat low but enough

12:40 - End of night

-- VincenzoTest - 11 Apr 2010
Topic revision: r10 - 21 Apr 2010, AdrianoFontana
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback